OpenStudy (anonymous):

Convert an equation from vertex form to Standard form?? f(x)= 2(x -11)^2 + 13

OpenStudy (anonymous):

@jim_thompson5910 @zepdrix @uri @Zarkon @ganeshie8 @radar @anonymous_user @triciaal

jimthompson5910 (jim_thompson5910):

are you able to expand out (x -11)^2 ?

OpenStudy (anonymous):

I'm not sure what that means. so far I've got: f(x)= 2(x -11)^2 + 13 \[(x - 11) (x - 11) \] \[x^2 - 11x -11x + 121\]

jimthompson5910 (jim_thompson5910):

-11x-11x turns into -22x

jimthompson5910 (jim_thompson5910):

so (x -11)^2 = x^2 - 22x + 121

jimthompson5910 (jim_thompson5910):

\[\Large f(x)= 2(x -11)^2 + 13\] \[\Large f(x)= 2(x^2 - 22x + 121) + 13\]

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

Distribute the 2?

jimthompson5910 (jim_thompson5910):

yep \[\Large f(x)= 2(x^2 - 22x + 121) + 13\] \[\Large f(x)= 2x^2 - 44x + 242 + 13\]

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

combine like terms.

OpenStudy (anonymous):

right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

giving you what final answer?

OpenStudy (anonymous):

\[f(x) = 2x^2 - 44x + 255 \]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct